I need to setup a quadratic expression and work it for this word problem. Could someone help me.
You have two rooms that are 12 x 24 that need wood flooring. Each box of wood flooring has 22.7 square ft in them. How many boxes would you have to have to finish both floors?

Answer by Theo(13342)

2 rooms are each 12 feet * 24 feet.

each box of flooring has 22.7 square feet in them.

the square feet in each room is 12 * 24 = 288

2 room has a total of 2 * 288 = 576 square feet.

divide this by the square feet in each box of flooring to get total number of boxes = 576 / 22.7 = 25.37444934.

round this to the next highest integer to get 26 boxes of flooring.

you will need approximately 26 boxes of flooring.

I don't see the need for a quadratic equation to solve this.

All you need to know is that the area of the floor in each room is equal to the length * the width.

You have the area of the 2 rooms.

You just divide that by the square feet in each box to get the number of boxes.